2 log 3x - log 9 = 1
Answer by stanbon(57410)

2 log(3x) - log 9 = 1
---
2[log(3)+log(x)] - 2log(3) = 1
2log(3) + 2log(x) - 2log(3) = 1
2log(x) = 1
log(x) = 1/2
x = 10^(1/2)
x = sqrt(10)
=====================
Cheers,
stan H.